some help here please, really confused.
The sum of three consecutive even integers is 72.
Write an equation you would use to find the integers.
x + (x + 1) + (x + 3) = 72
x + (x + 1) + (x + 2) = 72
4x = 72
x + (x + 2) + (x + 4) = 72

Found 2 solutions by Alan3354, unlockmath:
Answer by Alan3354(69443)

The sum of three consecutive even integers is 72.
Write an equation you would use to find the integers.
x + (x + 1) + (x + 3) = 72
x + (x + 1) + (x + 2) = 72
4x = 72
x + (x + 2) + (x + 4) = 72
------------------
The last one works.
x + (x + 2) + (x + 4) = 72
3x + 6 = 72
3x = 66
x = 22
--> 22, 24 & 26
----------------
I would do it the easy way, tho.
72/3 = 24 which is the middle number.
--> 22, 24 & 26

Answer by unlockmath(1688)

Hello,
The key to this is the sum of "EVEN" integers. For example 2+4+6 etc.
So to show this in an equation it would need to be set up as:
x+(x+2)+(x+4)=72
Another helpful hint: the word "is" in math problems usually mean "equal".
RJ
Check out a new book I wrote at:
www.math-unlock.com